OpenStudy (anonymous):

Not sure how to write this.... 3x (exponent of -2) ---------------- 6x (exponent of -5 How do I solve please.

OpenStudy (anonymous):

3/6=1/2 x^(-2)/x(-5)=x^(-2+5)=x^3 so, the answer is: 1/2 *x^3